find three consecutive even integers, such that 3 times the middle integer is 4 more than the sum of first and third
A.1,2,3
B.3,5,6
C.2,4,6
D.6,8,10





Answer by richwmiller(17219)

3*(y)=4+((x)+(z))
I set them up for you. Just multiply and add to see which equation is true. I did the first for you
Remember you looking for three consecutive even integers

A.1,2,3
3*2=4+(1+3)
6=8 is not true
also 1 and 3 are not even
B.3,5,6
3*5=4+(3+6)
C.2,4,6
3*4=4+(2+6)
D.6,8,10
3*(8)=4+(6+10)